# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

### v2

@@ Abstract
We propose semiparametric estimators called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models, relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical, which accommodates a skewness component. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations that prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows higher sparsity compared to the Gaussian copula graphical models.
Keywords: Elliptical distributions, robust statistics, skewness, and undirected graphical models.
@@ 1 Introduction
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
@@
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014) and in Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. However, as shown by Rossell and Zwiernik, (2021), the dependence structure of the elliptical distribution differs from that of the Gaussian case. In addition, a major pitfall of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
@@
We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators. These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in Arellano-Valle and Genton, 2010b . Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for a semiparametric estimation of graphical models that take into account distributional features such as directional skewness and heavy tails. The main idea is to consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions. The class of unified skew-elliptical distributions, which generalizes the closed skew-normal, enjoys key closure properties under marginalization, conditioning, and linear transformations. These properties enable e.g., to define conjugate priors in probit, tobit, and multinomial probit regression models based on $\operatorname{SU\mathcal{E}}$ distributions, as both the prior and posterior remain within the same family (Durante,, 2019, Karling et al.,, 2024).
The unified skew-elliptical distribution can be written in terms of the closed-skew normal, that possesses a stochastic representation defined through a normal distribution and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is then plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
The proposed estimators generalize the SKEPTIC estimator conforming to the meta skew-elliptical distribution and embed two new subclasses of estimators based on Kendall’s tau and Spearman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Gaussian data, the approach based on the meta skew-elliptical distribution allows us to quantify the asymmetry and the kurtosis presented in the data. The elliptical skew-(S)KEPTIC estimators possess skewness terms, encoded in the matrix of skewness parameters $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog,, 2002). The proposed framework extends to graphical models that incorporate both skewness and heavy tails, adding flexibility to nonparanormal-type data transformations.
@@
We prove that the elliptical skew-(S)KEPTIC estimators have deviation from the true correlation matrix $\Sigma^{0}$ that decreases as the sample size grows, showing a fast exponential rate of convergence in $\|\cdot\|_{\max}$ norm. The numerical analysis supports the theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimators prove an optimal graph recovery, showing statistical efficiency of the proposed method.
@@
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-normal distribution of Azzalini, (1985). They develop an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph. However, these studies do not focus on the direct estimation of the correlation matrix while incorporating skewness.
@@
The paper is organized as follows. In Section 2 we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution. In Section 3 we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators, whereas in Section 4 we provide some shrinkage procedures to estimate the precision matrix $\Omega$. Theoretical properties of the proposed estimator are discussed in Section 5. Section 6 presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns. Section 7 concludes the paper.
@@ 2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $\|A\|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
@@ 2.2 The Nonparanormal SKEPTIC
Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\hat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function. The transformation function $\hat{f}_{j}(x_{ij})$ is then defined as $\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function. In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\hat{F}_{j}(x_{ij})\approx\hat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$. Thus, the transformation results in
$\hat{f}_{j}(x_{ij})=\Phi^{-1}\left(\hat{r}_{ij}\right).$ (1)
@@
$\text{(Spearman's rho)}\ \hat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
$\text{(Kendall's tau)}\ \hat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
@@
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$. Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.

### v3

@@ Abstract
We propose semiparametric estimators, called elliptical skew-(S)KEPTIC, for efficiently and robustly estimating non-Gaussian graphical models. Our approach extends the semiparametric elliptical framework to the meta skew-elliptical family, which accommodates skewness. Theoretically, we show that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates for both graph recovery and parameter estimation. Through numerical simulations, we illustrate the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators. Finally, we apply the new method to the daily log-returns of the stocks in the S&P 500 index and obtain a sparser graph than with Gaussian copula graphical models.
Keywords: Elliptical distributions; Robust statistics; Skewness; Undirected graphical models.
@@ 1 Introduction
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in liu2009nonparanormal, which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. liu2012high propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
@@
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in han2014scale and in liu2012transelliptical, where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, liu2012transelliptical show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. A major drawback of the meta-elliptical class is that, while it includes skewed distributions, the dependence structure of its copula does not contain the skewness parameters, preventing them from playing a role in identifying the conditional independence structure.
@@
We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators. These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in arellano2010skew-e. Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for constructing graphical models by using a copula approach to estimate the second moment of the distribution taking into account the univariate directional skewness of the data. To do so, we consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in dominguez2003multivariate, which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions. The closed skew-normal distribution admits a stochastic representation involving both a normal and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The elliptical skew-(S)KEPTIC extends this estimator by incorporating a correction term that accounts for the directional skewness of the closed skew-normal distribution. Remarkably, we show that, given a bivariate relationship, concordant skewness directions strengthen the dependence between the variables, whereas discordant skewness directions weaken it.
The estimated correlation matrix is then plugged into existing shrinkage procedures, see, e.g., (meinshausen2006high; banerjee2008model; friedman2008sparse; yuan2010high; cai2011constrained). The elliptical skew-(S)KEPTIC estimators that generalize the SKEPTIC estimator are based on Kendall’s tau and Spearman’s rho statistics. When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (hult2002multivariate).
We prove that the elliptical skew-(S)KEPTIC estimators have the same rate of convergence of the SKEPTIC estimator of liu2012high plus a error rate from the estimate of the skewness parameters. The rates of convergence are studied in $\|\cdot\|_{\max}$ norm.
@@
We provide a numerical analysis to support the theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimators results in similar graph recovery as the SKEPTIC, which is proven to be optimal in liu2012high, showing statistical efficiency of the proposed method.
@@
Prior studies by zareifard2016skew, nghiem2022estimation, and sheng2023skewed have explored graphical models while incorporating skewness into the sparsity pattern estimation. zareifard2016skew exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach. nghiem2022estimation propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution. Instead, sheng2023skewed include the shape parameter in the determination of the precision matrix using the skew-normal distribution of azzalini1985class and developing an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph. However, these studies do not address the direct estimation of the correlation matrix. By contrast, we propose an estimator that explicitly incorporates the skewness parameters in the dependence structure through a copula-based approach.
@@
The paper is organized as follows. In Section 2, we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution. In Section 3, we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators whereas in Section 4, we provide some shrinkage procedures to estimate the precision matrix $\Omega$. Theoretical properties of the proposed estimator are discussed in Section 5. Section 6, presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns. Section 7 concludes the paper.
@@ 2 Background
In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of liu2012high, the semiparametric elliptical distributions in liu2012transelliptical and han2014scale. The section concludes with the introduction of the unified skew-elliptical family of distributions in arellano2010skew-e, and its link with the undirected graphical models.
@@ 2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $\|A\|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$, respectively, the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$, where $i\neq j$.
@@ 2.2 The Nonparanormal SKEPTIC
Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\widehat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function. The transformation function $\widehat{f}_{j}(x_{ij})$ is then defined as $\widehat{f}_{j}(x_{ij})=\Phi^{-1}(\widehat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function. In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\widehat{F}_{j}(x_{ij})\approx\widehat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$. Thus, the transformation results in:
$\widehat{f}_{j}(x_{ij})=\Phi^{-1}\left(\widehat{r}_{ij}\right).$ (1)
@@
$\text{(Spearman's rho)}\ \widehat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
$\text{(Kendall's tau)}\ \widehat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
Both $\widehat{\rho}_{ij}$ and $\widehat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to estimate the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$, that we recall.
@@
(kruskal1958ordinal, liu2012high): Suppose that there exists a set of $f=\left\{f_{1},\ldots,f_{p}\right\}$ monotonic functions such that $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, then we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\widehat{S}^{\rho}=\left[\widehat{S}^{\rho}_{ij}\right]$ and $\widehat{S}^{\tau}=\left[\widehat{S}^{\tau}_{ij}\right]$, where: $\widehat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\widehat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\widehat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\widehat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$. Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.